OpenStudy (natalie1234):

NEED HELP PLEASE!!! Multiply and simplify: 3xy^4z^2 • 5x^2yx

OpenStudy (chris911):

15x^4y^5

OpenStudy (chris911):

z^2

OpenStudy (campbell_st):

use the idea Numbers, Letters, Powers Signs.... so its 3*5xyz add the powers for each letter (the index law for multiplication) \[x^a \times x^b = x^{a+b}\] and since both terms are positve the answer is positive... last thing to know about powers a single letter is really \[x = x^1\]

OpenStudy (natalie1234):

okay thanks!(: